OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 19 feet, what is its lateral surface area? Use π = 22/7

OpenStudy (anonymous):

lateral surface area= h* 22/7*2*r

OpenStudy (anonymous):

base area = 22/7*2*r 154=22/7*2*r 154=44/7*r 154*7=44r (154*7)/44=r

OpenStudy (anonymous):

r = 24.5 LA = 19 * 22/7 * 2 * 24.5 LA = 2926? but it isn't one of my answer choices

OpenStudy (anonymous):

my answer choices are; 836 ft^2 990 ft^2 1672 ft^2 1980 ft^2 i tried solving it different ways before but it never ended up with the answer

OpenStudy (anonymous):

hold on

OpenStudy (accessdenied):

Base is a circle, so \(A_c = \pi r^2 \). We know that \(A_c = 154\). We want to find the circumference because the Lateral Surface Area is \(2 \pi r h\), the circumference of the circle times the height (since if we cut the cylinder down the side vertically and unrolled it, it'd become a rectangle with base as circumference and height stays the same). \[\begin{align} A_c = A_c \implies \pi r^2 &= 154\qquad \pi \approx \frac{22}{7} \\ \frac{22}{7}r^2 & = 154 \\ \end{align} \]

OpenStudy (anonymous):

so the radius will equal 7; which would fit, since 2 * 22/7 * 7 * 19 = 836 ft^2.

OpenStudy (accessdenied):

Yes. :)

OpenStudy (anonymous):

thanks alot , i was really stuck not getting what to do ☺

OpenStudy (accessdenied):

You're welcome. :) I had a few issues myself (I kept trying to use r=49, which led to strangely large results... lol)